OpenStudy (anonymous):

Write the sum or difference in simplest form. How much longer than 5/8 inch is 7/8 inch? A 1 1/2 B 1/4 C 1 3/8 D 1 1/8

OpenStudy (anonymous):

7/8-5/8=2/8=1/4 b

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }+\frac{ 7 }{ 8 }=\frac{ 12 }{ 8 }=\frac{ 3*4 }{ 2*4 }=\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }=\frac{ 2 }{ 2 }+\frac{ 1 }{ 2 }=1\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

it is asking the difference. so you subtract not add

OpenStudy (anonymous):

@koook is correct!(:

OpenStudy (anonymous):

(Write the SUM or difference in simplest form.)

OpenStudy (anonymous):

OR difference

OpenStudy (anonymous):

yes that is right .....

OpenStudy (anonymous):

The question is stating HOW MUCH LONGER 7/8inches is THAN 5/8 inches. So you would subtract!